### v2

@@ 3.2 Scene Structural Knowledge Transfer (SSKT)
After extracting structural knowledge from the ODI SR task, we propose the SSKT module with two components to transfer the extracted knowledge to the HR depth estimation task. Buttressed by the learned interpolation weights of CIIF, we share the parameters $g_{\theta}$ of CIIF between the two tasks. Besides, we propose the feature distillation (FD) loss to further utilize the scene structural knowledge contained in the features of the ODI SR task. The detail of the CIIF sharing strategy and FD loss is introduced in the following part.
@@
FD Loss: The CIIF sharing strategy can benefit the feature up-sampling process in the HR depth estimation task. To further enhance the structural details in the HR depth prediction, we propose the FD loss as a structural regularization to help two tasks learn common scene structural knowledge. Although ODIs and their corresponding depth maps are two different representations for the same scene with high geometric similarity, simply applying feature similarity loss, e.g., cosine similarity loss, introduces unnecessary textural knowledge and thus destroys the smoothness of the depth map. To address it, we propose to first transform the HR feature in the RGB part of ODI SR network to simulate the HR feature in the HR depth estimation network. Given an HR feature $F^{\mathrm{RGB}}_{\mathrm{HR}}$ as input, an encoder consisting of three convolutional layers and ReLU activation functions is applied to obtain a transformed feature $F^{\mathrm{T}}_{\mathrm{HR}}$. We then aim to maximize the cosine similarity between $F^{\mathrm{T}}_{\mathrm{HR}}$ and HR feature $F^{\mathrm{DE}}_{\mathrm{HR}}$ from the HR depth estimation network. FD loss can be formulated as follows:
@@
$\begin{split}\mathcal{L}_{\mathrm{T}}&=\mathcal{B}(\mathrm{M}(D^{\mathrm{T}}_{\mathrm{HR}}),D^{\mathrm{GT}}_{\mathrm{LR}}),\\ \mathcal{B}(x)&=\left\{\begin{aligned} &\left|x\right|,\left|x\right|\leq c\\ &\frac{x^{2}+c^{2}}{2c},\left|x\right|>c,\end{aligned}\right.\end{split}$ (8)
@@ 4.1 Implementation Details
Dataset. We conduct on three datasets: Stanford2D3D [29], Matterport3D [30] and 3D60 [14]. Stanford2D3D and Matterport3D are real-world datasets, and 3D60 is a synthetic dataset, provided by OmniDepth [14]. We follow UniFuse to split the three datasets respectively. The sizes of ODIs and paired depth maps are 512$\times$1024 in Stanford2D3D and Matterport3D, and 256$\times$512 in 3D60. On real-world datasets, the up-sampling factors are set to ×$\times 4$ and ×$\times 8$. We test on the synthetic dataset with ×$\times 8$ up-sampling factor to verify the generalization ability.
Implementation. We use three omnidirectional monocular depth estimation methods as backbones to extract features from the input ODIs: OmniDepth [14] with spherical convolutions, UniFuse-Fusion [10] with ERP type images and cube maps, and light-weight UniFuse-Equi [10] with only ERP type images. As for the feature extractor of the ODI SR task, we directly employ the commonly-used EDSR-baseline [54] to extract features from the input LR ODIs. We down-sample ODIs with Bicubic interpolation, while down-sample omnidirectional depth maps with Nearest interpolation. During training, we augment data with horizontal flipping and rolling. We collaboratively train the two tasks for 30 epochs using the Adam optimizer [55] with batch size 2. The initial learning rate is set to $41e-$ 4, and reduced by 2 times per 5 epochs. For hyper-parameters in $\mathcal{L}_{\mathrm{TOTAL}}$, we empirically set $\lambda_{1}=\lambda_{2}=10$, and $\lambda_{3}=\lambda_{4}=\lambda_{5}=1$. For fair comparison, we re-train the compared methods based on their provided settings.
@@ 4.2 Quantitative and Qualitative Evaluation
We evaluate our method with standard metrics, including mean absolute error (MAE), absolute relative error (Abs Rel), root mean square error in log space (RMSE log). We also utilize three accuracy measures ($\delta_{i}$), which are defined as the fraction of pixels where the relative error between the estimated depth and ground-truth depth is less than the threshold $i$, $i\in\{1.25,1.25^{2},1.25^{3}\}$. We only measure on valid pixels in the ground-truth depth map.
@@
Figure 5: The pipelines of baseline methods with five different settings. To be consistent with our weakly-supervised setting, these baseline methods only utilize LR ODIs as inputs and LR depth GT maps for supervision. MDE Net means the monocular depth estimation network, ISR Net means the ODI super-resolution network, Up Conv represents the learnable up-sampling convolutional layer, and Guide Up represents guided image filter [56].
@@
Fully-supervised methods. We compare with fully-supervised methods that utilize HR depth GT maps. To the best of our knowledge, there are no existing methods designed to generate HR depth prediction from an LR ODI. To this end, we add learnable up-sampling convolutional layers to OmniDepth [14], UniFuse-Fusion [10], and UniFuse-Equi [10] and re-train them with HR depth GT maps. As shown in Tab. 1, our weakly-supervised method not only achieves comparable results with corresponding fully-supervised methods (e.g., 3rd-4th rows), but also outperforms corresponding supervised methods in some cases (e.g., 1st-2nd rows). For example, in the 4th row of Fig. 4, our method predicts the bookshelf with regular shape, while the other corresponding fully-supervised methods predict the shape with more or less warp.
@@
The adaptability of our method in different input sizes is also verified, e.g., $128\times 256$ and $64\times 128$. Our weakly-supervised method outperforms the fully-supervised methods in several metrics, e.g., MAE, Abs Rel, and $\delta_{3}$. For example, with OmniDepth as the backbone with $128\times 256$ input size, our method obtains $3.8\%$ gain in MAE metric and $5.3\%$ gain in Abs Rel metric. When choosing a smaller input size $64\times 128$, our method can achieve similar results compared to the fully-supervised method, even though the LR depth GT map only contains 1/64 supervision signals of an HR depth GT map.
@@
In addition, we train on Matterport3D dataset with UniFuse-Fusion and OmniDepth as backbone, and achieve comparable results with the corresponding fully-supervised methods. For example, with UniFuse-Fusion as backbone and ×$\times 8$ up-sampling factor, our method improves $1.6\%$ in MAE metric and $0.3\%$ in $\delta_{3}$ metric. Finally, in Tab. 1, we verify the effectiveness of our method with higher resolution, i.e., $1024\times 2048$. Given the input resolution is $256\times 512$ and up-sampling factor is 4, the results show that our weakly-supervised method performs well. For example, compared with fully-supervised method, ours can improve 0.0050 in Abs Rel.
@@
1) Setting 1: As shown in Fig. 5(a), it is to estimate LR depth map from LR input ODI, and directly up-sample the LR depth prediction with Nearest interpolation. As shown in Fig. 6, our method outperforms baseline methods with Nearest interpolation in different input sizes and backbones. For example, with OmniDepth as the backbone, our method leads to 6.0% improvement in MAE metric, which proves the effects of structural knowledge transfer from the ODI SR task.
2) Setting 2: In Fig. 5(b), it is similar with Setting 1, except for replacing the Nearest interpolation with learnable up-sampling convolutional layers. The result of Setting 2 is 0.1830 Abs Rel metric, which is lower than our method (0.1585 Abs Rel). We analyze the reason is that the additional up-sampling convolutional layer can not be fully-supervised by HR depth GT maps, causing its performance dropped greatly.
3) Setting 3: In Fig. 5(c), it is to first up-sample LR ODIs with an ODI SR network, i.e., EDSR-baseline [54], which is fully-supervised with HR ground truth ODIs. Then, the super-resolved ODI is fed into a depth estimation network to generate an HR depth prediction. The HR depth prediction is down-sampled with Nearest interpolation to be compared with the LR depth GT map. In Tab. 2, the setting 3 performs badly, due to the super-resolved ODI contains much noise, which disturbs the following depth estimation network.
4) Setting 4: In Fig. 5(d), it is similar with Setting 3, except for that the depth estimation network is trained with LR ODI as input. We then directly transfer the pre-trained depth estimation network into the HR domain with the super-resolved ODI as input. However, it only obtains 0.4523 Abs Rel, which is worse than Setting 3 and our method. Its main drawback is that during training, the depth estimation network has no access to HR ODIs or HR depth maps.
@@
5) Setting 5: In Fig. 5(e), it first predicts a super-resolved image with an ODI SR network, i.e., EDSR-baseline [54], and an LR depth prediction with a depth estimation network. Then, it utilizes the super-resolved ODI to guide the interpolation process of the LR depth prediction with guided image filter [56] to obtain an HR depth prediction. However, the result is similar with that of Setting 4, due to the noise contained in the super-resolved image.
@@ 4.3 Ablation Studies
All the experiments are done on Stanford2D3D dataset with ×$\times 8$ up-sampling factor.
@@
Uncertainty estimation. In the first row of Tab. 3, we first omit uncertainty estimation in the ODI SR task, and share the parameters of CIIF between feature up-sampling processes in the ODI SR task and HR depth estimation task. Removing uncertainty estimation leads to 2.8% MAE and 7.2% Abs Rel reduction. Directly transferring knowledge between the ODI SR task and HR depth estimation task obtains sub-optimal performance, because the redundant textural details in the ODI SR task might interfere with the HR depth estimation task.
Sharing strategy of CIIF. To verify the effectiveness of our proposed sharing strategy, we additionally experiment on other combinations in Tab. 3. Specifically, in the second row, we choose to share the parameters of CIIF between feature up-sampling processes in the RGB part of the ODI SR task, and the HR depth estimation task. Moreover, we study the influence of sharing the parameters among the feature up-sampling processes in the RGB part of ODI SR task, the uncertainty part of ODI SR task, and the HR depth estimation task. However, we find that Abs Rel degrades 6.0% compared with sharing between the uncertainty part and HR depth estimation task. It also proves that textural details of directly transferring knowledge from features in the RGB part of ODI SR task can cause negative effects to the structural knowledge learning of the HR depth estimation task.
@@
The parameters of CIIF are all contained in an MLP. We share the parameters of CIIF between the ODI SR task and HR depth estimation task. To verify it, we conduct a momentum update [52] for the HR depth estimation task, which is considered as a asynchronous update. From Tab. 4, asynchronous update reduces performance greatly. We think the reason is that scene structural knowledge is essential for the HR depth estimation task, and asynchronous update destroys the common structural representations between the two tasks.
In addition, we investigate the effectiveness of cylindrical representation. As shown in Tab. 4, exploiting cylindrical representation obtains 5.1% gain in MAE metric compared with utilizing spherical angle coordinates. Note that using spherical angle coordinates is equal to using planar pixel coordinates in LIIF [44, 8]. The result demonstrates that cylindrical coordinate system is more proper to describe the spatial correlation of neighboring pixels in ERP type ODIs.
Loss function. By default, $\mathcal{L}_{\mathrm{UISR}}$ and $\mathcal{L}_{\mathrm{HD}}$ are necessary for the training. Therefore, we investigate the effectiveness of $\mathcal{L}_{\mathrm{LD}}$ and $\mathcal{L}_{\mathrm{FD}}$. In Tab. 5, without $\mathcal{L}_{\mathrm{LD}}$, the performance degrades 3.5% in Abs Rel metric. It demonstrates that $\mathcal{L}_{\mathrm{LD}}$ can utilize the LR depth GT maps to help obtain more accurate LR features for the HR depth estimation task. Without $\mathcal{L}_{FD}$, the performance also degrades 1.6% in RMSE log metric. In addition, we directly measure the similarity between the features in the ODI SR task without transform and features in the HR depth estimation task, and find that Abs Rel metric degrades from 0.1585 to 0.1723.
4.4 Discussion

### v3

@@ 3.2 Scene Structural Knowledge Transfer (SSKT)
After extracting structural knowledge from the ODI SR task, we propose the SSKT module with two components to transfer the extracted knowledge to the HR depth estimation task, as shown in Fig. 2. Buttressed by the learned interpolation weights of CIIF, we share the parameters $g_{\theta}$ of CIIF between the two tasks. This is because we think the learnable interpolation weights capture the relationships between pixels and thus represent the structural knowledge in both images and depth maps. Besides, to supplement high-resolution detailed features to the HR depth estimation task, we propose the feature distillation (FD) loss to transfer the scene structural knowledge from the ODI SR task, as shown in Fig. 4. The detail of the CIIF sharing strategy and FD loss is introduced in the following part.
@@
FD Loss: The CIIF sharing strategy can benefit the feature up-sampling process in the HR depth estimation task. To further enhance the structural details in the HR depth prediction, we propose the FD loss as a structural regularization to help two tasks learn common scene structural knowledge. Although ODIs and their corresponding depth maps are two different representations of the same scene with high geometric similarity, simply applying feature similarity loss, e.g., cosine similarity loss, introduces unnecessary textural knowledge and thus destroys the smoothness of the depth map. As shown in Fig. 4, we first transform the HR feature in the RGB part of the ODI SR network to simulate the HR feature in the HR depth estimation network. Given an HR feature $F^{\mathrm{RGB}}_{\mathrm{HR}}$ as input, an encoder consisting of three convolutional layers and ReLU activation functions is applied to obtain a transformed feature $F^{\mathrm{T}}_{\mathrm{HR}}$. We then aim to maximize the cosine similarity between $F^{\mathrm{T}}_{\mathrm{HR}}$ and HR feature $F^{\mathrm{DE}}_{\mathrm{HR}}$ from the HR depth estimation network. FD loss can be formulated as follows:
@@
$\begin{split}\mathcal{L}_{\mathrm{T}}&=\mathcal{B}(\mathrm{M}(D^{\mathrm{T}}_{\mathrm{HR}})-D^{\mathrm{GT}}_{\mathrm{LR}}),\\ \mathcal{B}(x)&=\left\{\begin{aligned} &\left|x\right|,\left|x\right|\leq c\\ &\frac{x^{2}+c^{2}}{2c},\left|x\right|>c,\end{aligned}\right.\end{split}$ (8)
@@ 4 Experiments
Figure 5: Visual comparison of fully-supervised method (i.e., UniFuse-Fusion) and ours on Matterport3D dataset. Best viewed in color.
@@ 4.1 Implementation Details
Dataset. We conduct on three datasets: Stanford2D3D [25], Matterport3D [26] and 3D60 [9]. Stanford2D3D and Matterport3D are real-world datasets, and 3D60 is a synthetic dataset, provided by OmniDepth [9]. We follow UniFuse to split the three datasets respectively. The sizes of ODIs and paired depth maps are 512$\times$1024 in Stanford2D3D and Matterport3D and 256$\times$512 in 3D60. On real-world datasets, the up-sampling factors are set to ×$\times 4$ and ×$\times 8$. We train, validate, and test on the two real-world datasets, which are more challenging and have higher resolution than the synthetic dataset. In addition, we test on the synthetic dataset with ×$\times 8$ up-sampling factor to verify the generalization ability.
Implementation. We use three omnidirectional depth estimation methods as backbones to extract features from the input ODIs: OmniDepth [9] with spherical convolutions, UniFuse-Fusion [5] with ERP type images and cube maps, and light-weight UniFuse-Equi [5] with only ERP type images. As for the feature extractor of the ODI SR task, we directly employ the commonly-used EDSR-baseline [41] to extract features from the input LR ODIs. We down-sample ODIs with Bicubic interpolation, while down-sample omnidirectional depth maps with Nearest interpolation. During training, we augment data with horizontal flipping and rolling. We collaboratively train the two tasks for 30 epochs using the Adam optimizer [42] with batch size 2. The initial learning rate is set to $41e-$ 4, and reduced by 2 times per 5 epochs. For hyper-parameters in $\mathcal{L}_{\mathrm{TOTAL}}$, we empirically set $\lambda_{1}=\lambda_{2}=10$, and $\lambda_{3}=\lambda_{4}=\lambda_{5}=1$. For a fair comparison, we re-train the compared methods based on their provided settings.
@@ 4.2 Quantitative and Qualitative Evaluation
We evaluate our method with standard metrics, including mean absolute error (MAE), absolute relative error (Abs Rel), and root mean square error in log space (RMSE log). We also utilize three accuracy measures ($\delta_{i}$), which are defined as the fraction of pixels where the relative error between the estimated depth and ground-truth depth is less than the threshold $i$, $i\in\{1.25,1.25^{2},1.25^{3}\}$. We only measure on valid pixels in the ground-truth depth map.
@@
Figure 6: The pipelines of baseline methods with five different settings. To be consistent with our weakly-supervised setting, these baseline methods only utilize LR ODIs as inputs and LR depth GT maps for supervision. MDE-Net means the monocular depth estimation network, ISR-Net means the ODI super-resolution network, Up-Conv represents the learnable up-sampling convolutional layer, and Guide-Up represents guided image filter [43].
@@
Fully-supervised methods. We compare with fully-supervised methods that utilize HR depth GT maps. To the best of our knowledge, there are no existing methods designed to generate HR depth prediction from an LR ODI. To this end, we add learnable up-sampling convolutional layers to OmniDepth [9], UniFuse-Fusion [5], and UniFuse-Equi [5] and re-train them with HR depth GT maps. As shown in Tab. 1, our weakly-supervised method not only achieves comparable results with corresponding fully-supervised methods (e.g., 3rd-4th rows) but also outperforms corresponding supervised methods in some cases (e.g., 1st-2nd rows). For example, in the 4th row of Fig. 5, our method predicts the furniture with regular shapes, while the other corresponding fully-supervised methods predict the shape with more or less warp.
@@
The adaptability of our method in different input sizes is also verified, e.g., $128\times 256$ and $64\times 128$. Our weakly-supervised method outperforms the fully-supervised methods in several metrics, e.g., MAE, Abs Rel, and $\delta_{3}$. For example, with OmniDepth as the backbone with $128\times 256$ input size, our method obtains $3.8\%$ gain in MAE metric and $5.3\%$ gain in Abs Rel metric. When choosing a smaller input size $64\times 128$, our method has inferior or comparable results compared with the fully-supervised methods. By taking ×$\times 8$ up-sampling factor and UniFuse-Fusion as the backbone, our method has no metric outperforming the fully supervised method on the Stanford2D3D dataset. We think it is because the LR depth GT map is too sparse, which only contains 1/64 supervision signals of an HR depth GT map.
@@
In addition, we train on the Matterport3D dataset with UniFuse-Fusion and OmniDepth as backbone and achieve comparable results with the corresponding fully-supervised methods. For example, with UniFuse-Fusion as the backbone and ×$\times 8$ up-sampling factor, our method improves $1.6\%$ in the MAE metric and $0.3\%$ in $\delta_{3}$ metric. Finally, in Tab. 1, we verify the effectiveness of our method with higher resolution, i.e., $1024\times 2048$. Given the input resolution is $256\times 512$ and the up-sampling factor is 4, the results show that our weakly-supervised method performs well. For example, compared with the fully-supervised method, ours can improve 0.0050 in Abs Rel.
@@
1) Setting 1: As shown in Fig. 6(a), it is to estimate LR depth map from LR input ODI, and directly up-sample the LR depth prediction with Nearest interpolation. As shown in Fig. 7, our method outperforms baseline methods with Nearest interpolation in different input sizes and backbones. For example, with OmniDepth as the backbone, our method leads to a 6.0% improvement in the MAE metric, which proves the effects of structural knowledge transfer from the ODI SR task.
2) Setting 2: In Fig. 6(b), it is similar to Setting 1, except for replacing the Nearest interpolation with learnable up-sampling convolutional layers. The result of Setting 2 is 0.1830 Abs Rel metric, which is lower than our method (0.1585 Abs Rel). We analyze the reason that the additional up-sampling convolutional layer can not be fully supervised by HR depth GT maps, causing its performance to drop greatly.
3) Setting 3: In Fig. 6(c), it is to first up-sample LR ODIs with an ODI SR network, i.e., EDSR-baseline [41], which is fully-supervised with HR ground truth ODIs. Then, the super-resolved ODI is fed into a depth estimation network to generate an HR depth prediction. The HR depth prediction is down-sampled with the Nearest interpolation to be compared with the LR depth GT map. In Tab. 2, setting 3 performs badly, due to the super-resolved ODI containing much noise, which disturbs the following depth estimation network.
4) Setting 4: In Fig. 6(d), it is similar to Setting 3, except that the depth estimation network is trained with LR ODI as input. We then directly transfer the pre-trained depth estimation network into the HR domain with the super-resolved ODI as input. However, it only obtains 0.4523 Abs Rel, which is worse than Setting 3 and our method. Its main drawback is that during training, the depth estimation network has no access to HR ODIs or HR depth maps.
@@
5) Setting 5: In Fig. 6(e), it first predicts a super-resolved image with an ODI SR network, i.e., EDSR-baseline [41], and an LR depth prediction with a depth estimation network. Then, it utilizes the super-resolved ODI to guide the interpolation process of the LR depth prediction with guided image filter [43]. However, the result is similar to that of Setting 4, due to the noise contained in the super-resolved image.
@@ 4.3 Ablation Studies
We conduct on the Stanford2D3D dataset with ×$\times 8$ factor.
@@
Uncertainty estimation. In the first row of Tab. 3, we first omit uncertainty estimation in the ODI SR task and share the parameters of CIIF between feature up-sampling processes in the ODI SR task and HR depth estimation task. Removing uncertainty estimation leads to 2.8% MAE and 7.2% Abs Rel reduction. Directly transferring knowledge between the ODI SR task and the HR depth estimation task obtains sub-optimal performance, because the redundant textural details in the ODI SR task might interfere with the HR depth estimation task.
Sharing strategy of CIIF. To verify the effectiveness of our proposed sharing strategy, we additionally experiment on other combinations in Tab. 3. Specifically, in the second row, we choose to share the parameters of CIIF between feature up-sampling processes in the RGB part of the ODI SR task, and the HR depth estimation task. Moreover, we study the influence of sharing the parameters among the feature up-sampling processes in the RGB part of the ODI SR task, the uncertainty part of the ODI SR task, and the HR depth estimation task. However, we find that Abs Rel degrades 6.0% compared with sharing between the uncertainty part and HR depth estimation task. It also proves that textural details of directly transferring knowledge from features in the RGB part of the ODI SR task can cause negative effects on the structural knowledge learning of the HR depth estimation task.
@@
The parameters of CIIF are all contained in an MLP. We share the parameters of CIIF between the ODI SR task and the HR depth estimation task. To verify it, we conduct a momentum update [39] for the HR depth estimation task, which is considered as an asynchronous update. From Tab. 4, asynchronous update reduces performance greatly. We think the reason is that scene structural knowledge is essential for the HR depth estimation task, and asynchronous update destroys the common structural representations between the two tasks.
In addition, we investigate the effectiveness of cylindrical representation. As shown in Tab. 4, exploiting cylindrical representation obtains a 5.1% gain in MAE metric compared with utilizing spherical angle coordinates. Note that using spherical angle coordinates is equal to using planar pixel coordinates in LIIF [31, 3]. The result demonstrates that the cylindrical coordinate system is more proper to describe the spatial correlation of neighboring pixels in ERP type ODIs.
@@
Table 6: Comparison of estimating uncertainty (Ours) with estimating the gradient for the ODI SR task.
MAE $\downarrow$ Abs Rel $\downarrow$ $\delta_{1}\uparrow$
Gradient 0.3141 0.1635 78.30
Uncertainty (Ours) 0.3135 0.1585 78.06
Loss function. By default, $\mathcal{L}_{\mathrm{UISR}}$ and $\mathcal{L}_{\mathrm{HD}}$ are necessary for the training. Therefore, we investigate the effectiveness of $\mathcal{L}_{\mathrm{LD}}$ and $\mathcal{L}_{\mathrm{FD}}$. In Tab. 5, without $\mathcal{L}_{\mathrm{LD}}$, the performance degrades 3.5% in Abs Rel metric. It demonstrates that $\mathcal{L}_{\mathrm{LD}}$ can utilize the LR depth GT maps to help obtain more accurate LR features for the HR depth estimation task. Without $\mathcal{L}_{FD}$, the performance also degrades 1.6% in the RMSE log metric. In addition, we omit the transformation process and find that the Abs Rel metric degrades from 0.1585 to 0.1723.
4.4 Discussion
Structural representation. To investigate the form of structural representation, we replace uncertainty estimation with gradient estimation [44]. In Tab. 6, uncertainty estimation performs better than gradient estimation. Compared with the gradient that only matches fixed edges of ODIs with the certain data distribution [24], uncertainty estimation is from the data itself and is more flexible for learning structural knowledge.